OpenStudy (anonymous):

What is the slope of the line (6, 9) and (-2, -3)? I think it's -3/4 but I'm not sure. Help please?

OpenStudy (anonymous):

Hmm, well crap then I got a lot of questions wrong if that's the right way!

hero (hero):

There's more than one way to solve math problems.

OpenStudy (anonymous):

y-9=m(x-6) y=mx-6m+9 -3=m(-2)-6m+9 -3=-8m+9 -8m=-12 m=3/2 hmmm we all got different answers lol

hero (hero):

y = mx + b 9 = 6m + b -3 = -2m + b 9 - 6m = b 2m - 3 = b b = b 9 - 6m = 2m - 3 9 + 3 = 2m + 6m 12 = 8m 12/8 = m 3/2 = m

OpenStudy (anonymous):

What I did was the whole m = y2 - y1 / x2 - x1

hero (hero):

\[m = \frac{9 -(-3)}{6 -(-2)}=\frac{9 + 3}{6 + 2} = \frac{12}{8} = \frac{3}{2}\]

hero (hero):

So there. Now you have have three ways to solve it.

OpenStudy (anonymous):

Hey check out this website it actually gives you the answer if you type in the points... thats how i got through geometry. :) http://www.webmath.com just click plots and geometry and fimnd whta you want under The equation of the line

OpenStudy (anonymous):

@Anders96 Haha I use that site sometimes :P N @Hero oohh! Yeah I was thinking that the negatives turn into positives, but I wasn't sure. Thanks guys!